OpenStudy (gabylovesyou):

The formula for degrees Celsius is: C = 5/9(F – 32), where F stands for degrees Fahrenheit. Part 1: Solve the formula for F. Show all steps. Part 2: Determine how many degrees Fahrenheit 6 degrees Celsius is.

OpenStudy (anonymous):

C = 5/9 (F - 32) 9/5 C = F-32 9/5 C + 32 = F